# Representations of categories of finite relational structures and associated endomorphism monoids

### 18 Mar 2026 v1

@@ 6.1. $\mathrm{CA}$} and $\mathrm{SA}$}
Suppose that $n\geqslant 2$ and $\lambda\in\operatorname{Irr}(G_{n})$ is singular. Then the projective cover $P^{c}_{n,\lambda}$ of $\Delta_{n,\lambda}$ has a filtration with factors $\Delta_{n,\lambda}$ and $\Delta_{n-1,\mu}$, where $(\mu,\lambda)$ is a pair appearing in Proposition 5.4.
@@
Since there is a surjective homomorphism $k{\mathscr{C}}\Psi_{n}\twoheadrightarrow\Delta_{n,\lambda}$, it follows that $P^{c}_{n,\lambda}$ is a direct summand of $k{\mathscr{C}}\Psi_{n}$. Consequently, it has a filtration with factors $\Delta_{n,\lambda}$ and some $\Delta_{n-1,\mu}$’s by Proposition 4.5. Furthermore, since $\Delta_{n,\lambda}$ has a simple top, we deduce that $P_{n,\lambda}^{c}$ is indecomposable. By Proposition 6.2, there is only a unique singular $\mu\in\operatorname{Irr}(G_{n-1})$ with multiplicity one such that $\Delta_{n-1,\mu}$ can occur in the filtration. ∎
We obtain the following result, which is a refined version of Theorem 4.7.
@@
The equivalence is given by the standard Morita equivalence with respect to the family of projective generators consisting of projective covers $P^{c}_{n,\lambda}$ of $\Delta_{n,\lambda}$ with $n\in\mathbb{Z}_{+}$ and $\lambda\in\operatorname{Irr}(G_{n})$. All statements concerning projectivity, irreducibility, and Loewy length have already been established in Propositions 5.5, 5.6, 6.2 and Corollary 6.3. Furthermore, if $\lambda\in\operatorname{Irr}(G_{n})$ is regular and $\mu\in\operatorname{Irr}(G_{m})$ is singular, then one has
@@ 6.2. $\mathrm{FA}$}
In this subsection let ${\mathscr{C}}={\mathrm{FA}}$ and $k$ be a field of characteristic 0. Note that the category $\mathscr{C}$$^{-}$ of finite sets and surjections is locally finite in the sense that for each object, there are only finitely many morphisms in $\mathscr{C}$$^{-}$ starting at it. Consequently, every representable functor $k{\mathscr{C}}^{-}(n,-)$ is of finite dimensional. Thus one can apply the algorithm in [28, Section 4] to construct the ordinary quiver of the category algebra $k{\mathscr{C}}^{-}$, whose vertices are pairs $(n,\lambda)$ with $n\in\mathbb{Z}_{+}$ and $\lambda\in\operatorname{Irr}(S_{n})$. Note that singular irreducible representations are sign representations by Proposition 5.4.
@@
We apply the quiver construction algorithm described in [28, Section 4]. In $\mathscr{C}$$^{-}$, the only unfactorizable morphisms with source $[n]$ are surjections from $[n]$ to [n1][n-1]. Up to the action of automorphism groups, all such surjections are equivalent, so we can choose the standard surjection $\alpha$ given by
@@
According to the algorithm in [28, Section 4], there exists an arrow from $(n,\lambda)$ to $(n-1,\mu)$ in the ordinary quiver if and only if $\lambda\downarrow_{G_{1}}^{G}$ and $\mu\downarrow_{H_{1}}^{H}$ share common simple summands appearing in the induced module
@@ 7. Uniformly continuous representations of endomorphism monoids
In this section, we establish a categorical equivalence between uniformly continuous modules over endomorphism monoids of infinite relational structures and sheaves on the corresponding finite substructure categories, and explore several concrete examples and applications.
@@ 7.1. Continuous and uniformly continuous representations
Let $(\Omega,\mathcal{R}_{G})$ be the relational structure determined by the action of a group $G$ on an infinite set $\Omega$ (here we do not require the action to be highly homogeneous), and let $\mathscr{E}\leqslant\operatorname{End}(\Omega,\mathcal{R}_{G})$ be a monoid consisting of some endomorphisms of this relational structure, equipped with the permutation topology. Specifically, we equip $\Omega$ with the discrete topology, $\Omega^{\Omega}$ with the product topology, and $\mathscr{E}$ with the subspace topology. Then any element $f\in\mathscr{E}$ has a basis of open neighborhoods described below:
@@
Note that in the above lemma, the finite substructure $A_{h,v}$ depends on both $v$ and $h$. In general, we cannot find a common $A_{v}$ working for all elements in $\mathscr{E}$. This stronger requirement is precisely the uniform continuity in topology, so we make the following definition.
@@
is continuous with respect to the permutation topology on all three components, but is not continuous when we equip the acting monoid $\mathscr{E}$ with the permutation topology and the other two $\mathscr{E}$ with the discrete topology. Indeed, for the identity $\operatorname{id}\in\mathscr{E}$ and a given $f\in\mathscr{E}$, we cannot find a finite substructure $A_{f}\subseteq\Omega$ such that when $g\in\mathscr{E}$ has the same restriction on $A$ as $f$, then $f\circ\operatorname{id}=g\circ\operatorname{id}$. Thus by Lemma 7.1, the action is not continuous.
The requirement of uniform continuity leads us to consider the submonoid of $\mathscr{E}$ consisting of endomorphisms whose image is finite. Explicitly, define
@@ 7.2. Realization and its right adjoint
Let $\mathscr{C}$} be the category whose objects are finite substructures of $(\Omega,\mathcal{R}_{G})$ and whose morphisms are restrictions of endomorphisms in $\mathscr{E}$, called the category of finite substructures associated to $\mathscr{E}$. Explicitly, $f:A\to B$ is a morphism in $\mathscr{C}$} if there is an endomorphism $\tilde{f}\in\mathscr{E}$ such that $\tilde{f}|_{A}=f$. For examples:
  • if $G=\operatorname{Sym}(\Omega)$ and $\mathscr{E}$ is the monoid of all maps from $\Omega$ to itself, then $\mathscr{C}$} is the category of finite subsets of $\Omega$ and maps;
  • if $G=\operatorname{Aut}(\Omega,\leqslant)$ with $\leqslant$ a dense linear order and $\mathscr{E}$ is the monoid of all order-preserving maps, then $\mathscr{C}$} is the category of finite subsets of $\Omega$ and order-preserving maps.
Note that $\mathscr{C}$} contains every inclusion morphism $\iota_{A,B}:A\hookrightarrow B$ if $A$ is a finite substructure of $B$.
@@
Let $\mathscr{C}$} be the category of finite substructures associated to $\mathscr{E}$, and $V$ a $\mathscr{C}$}-module. Then
@@
where $A$ ranges over all finite substructures of $\Omega$ (namely all finite subsets of $\Omega$ since each subset inherits a unique substructure from the global structure on $\Omega$), and the equivalence relation is generated by
@@
We check that this is a well-defined monoid action. Suppose $v]=[v^{\prime}]$ in $\mathbb{V}$, with $v\in V(A)$ and $v^{\prime}\in V(B)$. Then there exists a finite substructure $C$ such that $A,B\subseteq C$ and
@@
We complete the proof by showing the uniform continuity of this action. Equip $\mathbb{V}$ with the discrete topology. Let $v]\in\mathbb{V}$ be represented by $v\in V(A)$ for some finite substructure $A\subseteq\Omega$. For any $f,f^{\prime}\in\mathscr{E}$, if $f|_{A}=f^{\prime}|_{A}$, then by definition of the action,
@@
It is exact since the indexing category (finite substructures with inclusions) is filtered, and filtered colimits of $k$-modules are exact. Furthermore, it has a right adjoint satisfying some special properties.
@@
Definition of $\mathcal{L}$ on objects. Given a uniformly continuous $k\mathscr{E}$-module $M$, we construct a $\mathscr{C}$}-module $V=\mathcal{L}(M)$ as follows. For each finite substructure $A\subseteq\Omega$, define
@@
maps canonically to $M$ via $\phi:[v]\longmapsto v$. This map is surjective since every $v$ lies in some $V(A)$. To see that $\phi$ is well-defined, suppose $v]=[v^{\prime}]$ with $v\in V(A)$ and $v^{\prime}\in V(B)$. Then there exists a finite substructure $C\supseteq A,B$ such that
@@
Step 4: Adjunction. For $V\in{\mathscr{C}}{\textrm{-}\mathrm{Mod}}$ and $M\in k\mathscr{E}{\textrm{-}\mathrm{Mod}}^{\mathrm{uc}}$, define
@@ 7.3. Equivalences
By the classical result of Gabriel on localization of abelian categories [16, Proposition III.2.5] and Proposition 7.6, we have the following equivalence
@@
in other words, $J(A)$ is the set of all sieves each of which contains an inclusion $\iota_{A,B}:A\hookrightarrow B$ for a certain finite substructure $B$.
@@
is in $J(C)$, namely $f^{*}(S)$ contains an inclusion with source $C$. By the definition of $J(A)$, one can find an inclusion $\iota_{A,B}\in S$. Let $D$ be the finite substructure $B\cup C$. Note that the morphism $f:A\to C$ extends to an endomorphism $\tilde{f}\in\mathscr{E}$, whose restriction gives a morphism $B\to D$. Using the same argument as in the proof of Lemma 7.8, one can show $\iota_{C,D}\in f^{*}(S)$.
@@
The first equivalence is proved in Proposition 7.9, so we consider the second one. By [9, Corollary 1.2], it suffices to show that $J$-torsion elements defined in that paper are the same as torsion elements in Definition 7.7. To see this, given a $\mathscr{C}$}-module $V$ and an object $A$ in $\mathscr{C}$}, by [9, Definition 3.1], an element $v\in V(A)$ is $J$-torsion if and only if there is a covering sieve $S\in J(A)$ such that $f\cdot v=0$ for every $f\in S$. But $S$ contains an inclusion $\iota_{A,B}$, so $v$ is torsion in the sense of Definition 7.7. Conversely, if $v$ is torsion, then one can find an inclusion $\iota_{A,B}$ such that $\iota_{A,B}\cdot v=0$. Consequently, every morphism in the covering sieve $S_{A,B}\in J(A)$ sends $v$ to 0, so $v$ is $J$-torsion. ∎
@@ 7.4. Application I
In the first application, we take $\mathscr{E}=G\leqslant\operatorname{End}(\Omega,\mathcal{R}_{G})$. The associated category $\mathscr{C}$} is then the category of finite substructures and embeddings, called the age of $(\Omega,\mathcal{R}_{G})$; see [4]. Indeed, every morphism in $\mathscr{C}$} arises as the restriction of an element of $G$, and is therefore injective. Conversely, since $(\Omega,\mathcal{R}_{G})$ is homogeneous, every embedding extends to an element of $G$.
@@
If the action of $G$ on $\Omega$ satisfies the strong amalgamation property—namely, if the pointwise stabilizer of every finite substructure $A$ has no fixed point outside AA—then ${\mathscr{C}}^{{\mathrm{op}}}$ is isomorphic to the orbit category $\mathscr{G}$ of $G$ with respect to the cofinal system of stabilizer subgroups of finite substructures; see [29, Theorem 1.4]. In this case, Theorem 7.11 yields an equivalence
@@
If, on the other hand, $\mathscr{C}$} fails to satisfy the strong amalgamation property, then there exist non-isomorphic finite substructures $A$ and $B$ with identical pointwise stabilizer subgroups, $G_{A}=G_{B}$. In this situation, the opposite category ${\mathscr{C}}^{{\mathrm{op}}}$ is no longer isomorphic to the orbit category $\mathscr{G}$: the former distinguishes $A$ and $B$ as separate objects, while the latter identifies them. Nevertheless, this distinction disappears upon sheafification. Although ${\mathscr{C}}^{{\mathrm{op}}}$ and $\mathscr{G}$ may differ as sites, they give rise to equivalent sheaf categories. Conceptually, the atomic topology forces finite substructures with the same stabilizer subgroup to become locally indistinguishable, so that the resulting topos depends only on the stabilizer data rather than on the individual finite substructures. More formally, one may apply the Comparison Lemma to obtain an equivalence

### 25 Apr 2026 v3

@@ 6.1. $\mathrm{CA}$} and $\mathrm{SA}$}
Let $k$ be a field and $\lambda\in\operatorname{Irr}(G_{n})$. One has:
  1. (1)
    if $\lambda$ is regular, then $\Delta_{n,\lambda}$ is projective;
  2. (2)
    if $n=1$, then $\Delta_{n,\lambda}$ is projective of length 2;
  3. (3)
    if $n\geqslant 2$ and $\lambda$ is singular, then the projective cover $P^{c}_{n,\lambda}$ of $\Delta_{n,\lambda}$ has a filtration with factors $\Delta_{n,\lambda}$ and $\Delta_{n-1,\mu}$, where $(\mu,\lambda)$ is a pair appearing in Proposition 5.4.
@@
Since there is a surjective homomorphism $k{\mathscr{C}}\Psi_{n}\twoheadrightarrow\Delta_{n,\lambda}$, it follows that $P^{c}_{n,\lambda}$ is a direct summand of $k{\mathscr{C}}\Psi_{n}$. Consequently, it has a filtration with factors $\Delta_{n,\lambda}$ and some $\Delta_{n-1,\mu}$’s by Proposition 4.5. Furthermore, since $\Delta_{n,\lambda}$ has a simple top, we deduce that $P_{n,\lambda}^{c}$ is indecomposable.
(1) By Proposition 5.3, $P_{n,\lambda}$ has a filtration by $\Delta_{n,\lambda}$ and some $\Delta_{i,\mu}$ with $i<n$. But Proposition 6.2 tells us that there is no nontrivial extension of $\Delta_{n,\lambda}$ by other indecomposable standard modules. Consequently, $\Delta_{n,\lambda}$ is a direct summand of $P_{n,\lambda}$, and hence is projective.
(2) This is clear since $\Delta_{1,\lambda}\cong k{\mathscr{C}}(1,-)$.
(3) By Proposition 6.2, there is only a unique singular $\mu\in\operatorname{Irr}(G_{n-1})$ with multiplicity one such that $\Delta_{n-1,\mu}$ can occur in the filtration. ∎
If $k$ has characteristic 0, we obtain the following result, a refined version of Theorem 4.7.
@@
The equivalence is given by the standard Morita equivalence with respect to the family of projective generators consisting of projective covers $P^{c}_{n,\lambda}$ of $\Delta_{n,\lambda}$ with $n\in\mathbb{Z}_{+}$ and $\lambda\in\operatorname{Irr}(G_{n})$. All statements concerning projectivity, irreducibility, and Loewy length have already been established in Propositions 5.6, 5.9, and Corollary 6.3. Furthermore, if $\lambda\in\operatorname{Irr}(G_{n})$ is regular and $\mu\in\operatorname{Irr}(G_{m})$ is singular, then one has
@@ 6.2. $\mathrm{FA}$}
In this subsection let ${\mathscr{C}}={\mathrm{FA}}$ and $k$ be a field of characteristic 0. Note that the category $\mathscr{C}$$^{-}$ of finite sets and surjections is locally finite in the sense that for each object, there are only finitely many morphisms in $\mathscr{C}$$^{-}$ starting at it. Consequently, every representable functor $k{\mathscr{C}}^{-}(n,-)$ is of finite dimensional. Thus one can apply the algorithm in [30, Section 4] to construct the ordinary quiver of the category algebra $k{\mathscr{C}}^{-}$, whose vertices are pairs $(n,\lambda)$ with $n\in\mathbb{Z}_{+}$ and $\lambda\in\operatorname{Irr}(S_{n})$. Note that singular irreducible representations are sign representations by Proposition 5.4.
@@
We apply the quiver construction algorithm described in [30, Section 4]. In $\mathscr{C}$$^{-}$, the only unfactorizable morphisms with source $[n]$ are surjections from $[n]$ to [n1][n-1]. Up to the action of automorphism groups, all such surjections are equivalent, so we can choose the standard surjection $\alpha$ given by
@@
According to the algorithm in [30, Section 4], there exists an arrow from $(n,\lambda)$ to $(n-1,\mu)$ in the ordinary quiver if and only if $\lambda\downarrow_{G_{1}}^{G}$ and $\mu\downarrow_{H_{1}}^{H}$ share common simple summands appearing in the induced module
@@ 7. Uniformly continuous representations of endomorphism monoids
In this section, we establish a categorical equivalence between uniformly continuous modules over topological transformation monoids $\mathscr{E}$ and sheaves on the corresponding categories $\mathscr{C}$} of finite subsets and finite restrictions of elements in $\mathscr{E}$, and explore several concrete examples and applications.
@@ 7.1. Continuous and uniformly continuous representations
Let $\Omega$ be an infinite set, and let $\mathscr{E}\leqslant\Omega^{\Omega}$ be a monoid of self-maps equipped with the permutation topology. Specifically, we equip $\Omega$ with the discrete topology, $\Omega^{\Omega}$ with the product topology, and $\mathscr{E}$ with the subspace topology. Then any element $f\in\mathscr{E}$ has a basis of open neighborhoods described below:
@@
Note that in the above lemma, the finite subset $A_{h,v}$ depends on both $v$ and $h$. In general, we cannot find a common $A_{v}$ working for all elements in $\mathscr{E}$. This stronger requirement is precisely the uniform continuity in topology, so we make the following definition.
@@
is continuous with respect to the permutation topology on all three components, but is not continuous when we equip the acting monoid $\mathscr{E}$ with the permutation topology and the other two $\mathscr{E}$ with the discrete topology. Indeed, for the identity $\operatorname{id}\in\mathscr{E}$ and a given $f\in\mathscr{E}$, we cannot find a finite subset $A_{f}\subseteq\Omega$ such that when $g\in\mathscr{E}$ has the same restriction on $A$ as $f$, then $f\circ\operatorname{id}=g\circ\operatorname{id}$. Thus by Lemma 7.1, the action is not continuous.
The requirement of uniform continuity leads us to consider the submonoid of $\mathscr{E}$ consisting of maps whose image is finite. Explicitly, define
@@ 7.2. Realization and its right adjoint
Given a transformation monoid $\mathscr{E}\leqslant\Omega^{\Omega}$, we construct a category $\mathscr{C}$} as follows. Objects in $\mathscr{C}$} are finite subsets of $\Omega$, and morphisms are restrictions of elements of $\mathscr{E}$. Call $\mathscr{C}$} the category of finite subset associated to $\mathscr{E}$. Explicitly, $f:A\to B$ is a morphism in $\mathscr{C}$} if there is an element $\tilde{f}\in\mathscr{E}$ such that $\tilde{f}|_{A}=f$. For examples:
  • if $\mathscr{E}=\Omega^{\Omega}$, then $\mathscr{C}$} is the category of finite subsets of $\Omega$ and maps;
  • if $\Omega$ is equipped with a linear order $\leqslant$ and $\mathscr{E}=\operatorname{End}(\Omega,\leqslant)$, then $\mathscr{C}$} is the category of finite subsets of $\Omega$ and order-preserving maps.
Note that $\mathscr{C}$} contains every inclusion morphism $\iota_{A,B}:A\hookrightarrow B$ if $A$ is a finite subset of $B$.
@@
Let $\mathscr{C}$} be the category of finite subset associated to $\mathscr{E}$, and $V$ a $\mathscr{C}$}-module. Then
@@
where $A$ ranges over all finite subsets of $\Omega$, and the equivalence relation is generated by
@@
We check that this is a well-defined monoid action. Suppose $v]=[v^{\prime}]$ in $\mathbb{V}$, with $v\in V(A)$ and $v^{\prime}\in V(B)$. Then there exists a finite subset $C$ such that $A,B\subseteq C$ and
@@
We complete the proof by showing the uniform continuity of this action. Equip $\mathbb{V}$ with the discrete topology. Let $v]\in\mathbb{V}$ be represented by $v\in V(A)$ for some finite subset $A\subseteq\Omega$. For any $f,f^{\prime}\in\mathscr{E}$, if $f|_{A}=f^{\prime}|_{A}$, then by definition of the action,
@@
It is exact since the indexing category (finite subsets with inclusions) is filtered, and filtered colimits of $k$-modules are exact. Furthermore, it has a right adjoint satisfying some special properties.
@@
Definition of $\mathcal{L}$ on objects. Given a uniformly continuous $k\mathscr{E}$-module $M$, we construct a $\mathscr{C}$}-module $V=\mathcal{L}(M)$ as follows. For each finite subset $A\subseteq\Omega$, define
@@
maps canonically to $M$ via $\phi:[v]\longmapsto v$. This map is surjective since every $v$ lies in some $V(A)$. To see that $\phi$ is well-defined, suppose $v]=[v^{\prime}]$ with $v\in V(A)$ and $v^{\prime}\in V(B)$. Then there exists a finite subset $C\supseteq A,B$ such that
@@
Adjunction. For $V\in{\mathscr{C}}{\textrm{-}\mathrm{Mod}}$ and $M\in k\mathscr{E}{\textrm{-}\mathrm{Mod}}^{\mathrm{uc}}$, define
@@ 7.3. Equivalences
By the classical result of Gabriel on localization of abelian categories [18, Proposition III.2.5] and Proposition 7.6, we have the following equivalence
@@
in other words, $J(A)$ is the set of all sieves each of which contains an inclusion $\iota_{A,B}:A\hookrightarrow B$ for a certain finite subset $B$.
@@
is in $J(C)$, namely $f^{*}(S)$ contains an inclusion with source $C$. By the definition of $J(A)$, one can find an inclusion $\iota_{A,B}\in S$. Let $D$ be the finite subset $B\cup C$. Note that the morphism $f:A\to C$ extends to a map $\tilde{f}\in\mathscr{E}$, whose restriction gives a morphism $B\to D$. Using the same argument as in the proof of Lemma 7.8, one can show $\iota_{C,D}\in f^{*}(S)$.
@@
The first equivalence is proved in Proposition 7.9, so we consider the second one. By [11, Corollary 1.2], it suffices to show that $J$-torsion elements defined in that paper are the same as torsion elements in Definition 7.7. To see this, given a $\mathscr{C}$}-module $V$ and an object $A$ in $\mathscr{C}$}, by [11, Definition 3.1], an element $v\in V(A)$ is $J$-torsion if and only if there is a covering sieve $S\in J(A)$ such that $f\cdot v=0$ for every $f\in S$. But $S$ contains an inclusion $\iota_{A,B}$, so $v$ is torsion in the sense of Definition 7.7. Conversely, if $v$ is torsion, then one can find an inclusion $\iota_{A,B}$ such that $\iota_{A,B}\cdot v=0$. Consequently, every morphism in the covering sieve $S_{A,B}\in J(A)$ sends $v$ to 0, so $v$ is $J$-torsion. ∎
@@ 7.4. Application I
In the first application, we take $\mathscr{E}=G\leqslant\operatorname{Sym}(\Omega)$ to be a permutation group on $\Omega$. The associated category $\mathscr{C}$} has finite subsets of $\Omega$ as objects, and morphisms given by restrictions of elements of $G$. Equivalently, $\mathscr{C}$} can be identified with the category of finite substructures and embeddings of $(\Omega,\mathcal{R}_{G})$, where $\mathcal{R}_{G}$ is the orbit relational structure induced by the action of $G$ on $\Omega$; see [4] for details. Every morphism in $\mathscr{C}$} is injective, as it arises from a restriction of a permutation in $G$. Conversely, since $(\Omega,\mathcal{R}_{G})$ is homogeneous, every embedding of finite substructures extends to an element of $G$.
@@
If the action of $G$ on $\Omega$ satisfies the strong amalgamation property—namely, if the pointwise stabilizer of every finite subset $A$ has no fixed point outside AA—then ${\mathscr{C}}^{{\mathrm{op}}}$ is isomorphic to the orbit category $\mathscr{G}$ of $G$ with respect to the cofinal system of stabilizer subgroups of finite subsets; see [31, Theorem 1.4]. In this case, Theorem 7.11 yields an equivalence
@@
If, on the other hand, $\mathscr{C}$} fails to satisfy the strong amalgamation property, then there exist non-isomorphic finite subsets $A$ and $B$ with identical pointwise stabilizer subgroups, $G_{A}=G_{B}$. In this situation, the opposite category ${\mathscr{C}}^{{\mathrm{op}}}$ is no longer isomorphic to the orbit category $\mathscr{G}$: the former distinguishes $A$ and $B$ as separate objects, while the latter identifies them. Nevertheless, this distinction disappears upon sheafification. Although ${\mathscr{C}}^{{\mathrm{op}}}$ and $\mathscr{G}$ may differ as sites, they give rise to equivalent sheaf categories. Conceptually, the atomic topology forces finite subsets with the same stabilizer subgroup to become locally indistinguishable, so that the resulting topos depends only on the stabilizer data rather than on the individual finite subsets. More formally, one may apply the Comparison Lemma to obtain an equivalence